OpenStudy (anonymous):

Show that if the vectors v, w are orthogonal then ||v+w||^2 = ||v||^2 + ||w||^2

OpenStudy (anonymous):

\[\Large \vec{v}\cdot \vec{v}=vv\cos0=v^2 \] so \[\Large(\vec{v}+\vec{w})\cdot(\vec{v}+\vec{w})=|| \vec{v}+\vec{w}||=\vec{v}\cdot\vec{v}+2\vec{v}\cdot \vec{w}+\vec{w}\cdot\vec{w} \]

OpenStudy (anonymous):

Know how to complete from here?

OpenStudy (anonymous):

No I'm completely lost

OpenStudy (anonymous):

Well I only applied the definition of length through the scalar product and expanded it from one vector to two vectors (which happen to be orthogonal to one another), orthogonal means that: \[\Large \vec{v}\cdot \vec{w}=0 \] As by the scalar product: \[\Large \vec{a}\cdot \vec{b}=||\vec{a}|| \cdot ||\vec{b}|| \cos\gamma \] for gamma=90 degree, the right hand side becomes zero, therefore orthogonal vectors must turn out to be zero when scalar multiplied.